# General Hamiltonian description of nonreciprocal interactions

### v3

@@ Appendix A Constrained Hamiltonian embedding of nonreciprocal systems
Figure 5: Schematic representation of various models exhibiting nonreciprocal equations of motion. (a) Metal-dielectric self-propelled Janus colloidal particles (Sec. A.2) experience self-propulsion, reciprocal repulsive forces and nonreciprocal torques. (b) Self-aligning active particles (Sec. A.3) interact via reciprocal repulsive forces and exhibit self-propulsion. The nonreciprocal torques cause each particle to rotate toward its velocity direction. (c) Programmed robotic metamaterials (Sec. A.4) consist of robotic units connected by springs that exert asymmetric forces on the left and right sides. (d) Feedback-controlled active particles (Sec. A.5) interact via reciprocal forces and self-propel when a sufficient number of neighboring particles enter their vision cone.
A.1 XY spins with vision cone interactions
A first illustrative example is the vision-cone XY spins, $\mathbf{S}_{i}=(\cos\theta_{i},\sin\theta_{i})$, on a square lattice discussed in the Main Text. Here, the nonreciprocal equations of motion take the form [cf. Eq. (1)]:
@@ A.1 XY spins with vision cone interactions
The equations of motion are given in Eq. (4), ignoring the noise term. The origin and technical handling of the constraint are discussed in detail in Sec. B.
@@
A.2 Metal-dielectric Janus particles
@@
To explain these phenomena, previous work proposed a two-dimensional microscopic model based on the dipolar interactions between the hemispheres of the particles (Fig. 5a) [15]. The interaction torques, which tend to rotate a particle either away or towards the position of a second particle, are intrinsically nonreciprocal; they do not have a reciprocal limit [15, 20]. The position and orientation of particle $i$ are denoted by $\mathbf{r}_{i}$ and $\mathbf{\hat{n}}_{i}$, respectively. The equations of motion are given by
@@
which is the same as Eq. (A.2), except for the noise terms $\bm{\eta}_{i}^{\text{t}},\bm{\eta}_{i}^{\text{r}}$.
A.3 Self-aligning active particles
@@
Here, $v_{0}$ denotes the self-propulsion speed, $\gamma_{\text{t}}$ is the translational friction coefficient, and $\mathbf{F}_{ji}=-\mathbf{F}_{ij}$ represents the reciprocal repulsive force that particle $i$ feels from particle $j$. Such interaction force can be derived from a potential $U(|\mathbf{r}_{i}-\mathbf{r}_{j}|)$ between particles $i$ and $j$. The second equation expresses that the orientation $\hat{\mathbf{n}}_{i}$ of a particle rotates towards its velocity $\dot{\mathbf{r}}_{i}$ over an alignment length $\ell_{\text{a}}$. A schematic illustration is presented in Fig. 5b. Ignoring the noise terms $\bm{\eta}_{i}^{\text{t}}$ and $\bm{\eta}_{i}^{\text{r}}$, we substitute the velocity in the second equation using the first equation, thereby obtaining the following expression for the equations of motion
@@
which is equivalent to Eq. (A.3), except for the noise terms.
A.4 Programmed robotic metamaterials
Let us now look at mechanical metamaterials made of robots whose interactions are programmed to break reciprocity  [23, 24]. In such materials, nonreciprocal interactions give rise to spatially asymmetric standing waves and unidirectionally amplified propagating waves, realizing the mechanical analogue of the non-Hermitian skin effect [73, 74]. Specifically, the robots are programmed to turn more (less) depending on how their left (right) neighbors turn, as schematized in Fig. 5c. The dynamics can be mapped onto the following nonreciprocal discrete wave equation, which arises from a mass-spring model in which the forces exerted by the springs on the left and right sides differ [23]:
@@ A.4 Programmed robotic metamaterials
Here $u_{i}$ denotes the displacement of robot $i$; $k\left(1+\varepsilon\right)$ and $k\left(1-\varepsilon\right)$ represent the stiffnesses of coupling springs from left to right and from right to left, respectively, with the dimensionless quantity $\varepsilon$ quantifying non-reciprocity. For convenience, we have set the mass to unity.
@@
We observe that the kinetic energy of the auxiliary particles is negative. Such a negative kinetic energy term in the auxiliary system is necessary to enforce our constraints; a more detailed explanation will be provided in Sec. B.2.
@@
A.5 Feedback-controlled active particles with vision-cone interactions
@@
where $V_{i}^{\alpha}$ denotes the vision cone of angle $\alpha$ around the orientation of particle $i$ (Fig. 5d). Respectively, $r_{ij}$ is the distance between particles $i$ and $j$, so that the perception metric assigns less weight to more distant particles. If a given particle has a perception metric $P_{i}>P^{*}$, it is made to propel at speed $v_{0}$, which is achieved by shining light on it through the laser scanning system. Otherwise, the particle remains passive. Thus, particle motion can be modeled as
@@
which is the same as Eq. (A.5), except for the noise terms.
Appendix B Symplectic structure of the Hamiltonian embedding: Implementing the constraint for interactions with overdamped and undamped nonreciprocal dynamics
In Sec. A, we introduced Hamiltonian embeddings and outlined the appropriate constraints for several nonreciprocal systems, both for damped and undamped dynamics. Here, we elaborate on these two cases separately, discussing in detail the implementation of the constraint and the corresponding symplectic structure. Although the discussion below is general, for concreteness we shall illustrate the procedure using the vision-cone XY spins as an example, for which the interaction $J_{ij}(\theta_{i})$ takes the form
@@
B.1 Overdamped nonreciprocal systems
@@
Using Eq. (B.1), the equation of motion of the constraint is given by
@@
$\sin\left(\theta_{i}-\theta_{j}\right)+\sin\left(\theta_{i}-\varphi_{j}\right)$ $=$ $0$ (57)
@@
B.2 Undamped nonreciprocal systems
@@
Using Eq. (B.2), the equations of motion of the constraints, $G^{(1)}_{i}$ and $G^{(2)}_{i}$, are given by
@@
$\sin\left(\theta_{i}-\theta_{j}\right)+\sin\left(\varphi_{i}-\theta_{j}\right)$ $=$ $0$ (68)
@@
The occurrence of negative kinetic energy for the auxiliary degrees of freedom can be explained as follows: Consider first the free system where the degrees of freedom are noninteracting, $\mathcal{U}=0$, and both terms in the kinetic energy are positive; then, by conservation of angular momentum $L^{\varphi}_{j}+L^{\theta}_{j}=0$, the $\theta$ and $\varphi$ degrees of freedom will rotate opposite to one another, thus violating the constraint that pins them to move together (i.e., in the same direction). To enforce the latter without over-constraining the dynamics of the resulting joint motion, a negative kinetic energy is required. Returning to the Hamiltonian (B.2), more formally, as discussed above, we need to construct a constraint as $\theta_{i}(t)=\varphi_{i}(t)+\pi$. Taking two time derivatives, we have $\ddot{\theta}_{i}(t)=\ddot{\varphi}_{i}(t)$. If the interactions between the original and the auxiliary degrees of freedom are defined as in Eq. (21), the torques experienced by the original spin and the auxiliary spin will be oppositely directed under the imposed constraint, as given by $-\sum_{j\in\langle ij\rangle}J_{ij}\sin(\theta_{i}-\theta_{j})=-\big{[}-\sum_{j\in\langle ij\rangle}J_{ij}\sin(\varphi_{i}-\theta_{j})\big{]}$. To maintain the opposite sign in the EOM of the angular momenta, the kinetic energy for the auxiliary degree of freedom has a negative sign, $-\frac{1}{2}\dot{\varphi}_{i}^{2}$. Such negative kinetic energy terms for the auxiliary degree of freedom have been discussed in Ref. [56] where they arise as a result of reformulating the action principle as an initial- (rather than a boundary-) value problem. Similarly, in the programmed robotic metamaterials example from Sec. A.4, we used the kinetic term $\frac{1}{2}\sum_{i}\left(p_{i}^{u}\right)^{2}-\left(p_{i}^{w}\right)^{2}$ to construct the Hamiltonian. Such anisotropic kinetic energy terms have also been encountered before in the description of azimuthal curl forces [76].
@@
Appendix C Equivalence between nonreciprocal Langevin dynamics and Glauber Monte Carlo dynamics
This section demonstrates that the steady-state density distributions resulting from the original nonreciprocal Langevin dynamics and from the constrained Glauber dynamics based on the Hamiltonian that we propose are identical. In Sec. C.1, we first derive a Fokker-Planck-like equation for the phase-space density distribution obtained from Langevin dynamics; then, in Sec. C.2, we prove that it is equivalent to the corresponding distribution obtained using constrained Glauber dynamics.
@@ Appendix C Equivalence between nonreciprocal Langevin dynamics and Glauber Monte Carlo dynamics
although the procedure is general and applies to other models (see Sec. A).
C.1 Langevin steady-state distribution

### v4

@@ Appendix B Examples of nonreciprocal systems treated within the Hamiltonian embedding
Figure 6: Schematic representation of various models exhibiting nonreciprocal equations of motion. (a) XY spins with vision-cone interactions (Sec. B.1) experiences torques from the neighboring spins within its vision cone. (b) Metal-dielectric self-propelled Janus colloidal particles (Sec. B.2) experience self-propulsion, reciprocal repulsive forces and nonreciprocal torques. (c) Self-aligning active particles (Sec. B.3) interact via reciprocal repulsive forces and exhibit self-propulsion. The nonreciprocal torques cause each particle to rotate toward its velocity direction. (d) Programmed robotic metamaterials (Sec. B.4) consist of robotic units connected by springs that exert asymmetric forces on the left and right sides. (e) Feedback-controlled active particles (Sec. B.5) interact via reciprocal forces and self-propel when a sufficient number of neighboring particles enter their vision cone. (f) Sedimenting particles settling through a viscous fluid under gravity (Sec. B.6) affect each other by generating flow fields. The resulting interactions are intrinsically nonreciprocal [61, 38].
B.1 XY spins with vision cone interactions
A first illustrative example is the vision-cone XY spins, $\mathbf{S}_{i}=(\cos\theta_{i},\sin\theta_{i})$, on a square lattice discussed in the Main Text. Here, the nonreciprocal equations of motion take the form [cf. Eq. (1) and Fig. 6(a)]:
@@ B.1 XY spins with vision cone interactions
The equations of motion are given in Eq. (4), ignoring the noise term. The origin and technical handling of the constraint are discussed in detail in Sec. C.
@@
B.2 Metal-dielectric Janus particles
@@
To explain these phenomena, previous work proposed a two-dimensional microscopic model based on the dipolar interactions between the hemispheres of the particles (Fig. 6b) [89]. The interaction torques, which tend to rotate a particle either away or towards the position of a second particle, are intrinsically nonreciprocal; they do not have a reciprocal limit [89, 21]. The position and orientation of particle $i$ are denoted by $\mathbf{r}_{i}$ and $\mathbf{\hat{n}}_{i}$, respectively. The equations of motion are given by
@@
which is the same as Eq. (B.2), except for the noise terms $\bm{\eta}_{i}^{\text{t}},\bm{\eta}_{i}^{\text{r}}$.
B.3 Self-aligning active particles
@@
Here, $v_{0}$ denotes the self-propulsion speed, $\gamma_{\text{t}}$ is the translational friction coefficient, and $\mathbf{F}_{ji}=-\mathbf{F}_{ij}$ represents the reciprocal repulsive force that particle $i$ feels from particle $j$. Such interaction force can be derived from a potential $U(|\mathbf{r}_{i}-\mathbf{r}_{j}|)$ between particles $i$ and $j$. The second equation expresses that the orientation $\hat{\mathbf{n}}_{i}$ of a particle rotates towards its velocity $\dot{\mathbf{r}}_{i}$ over an alignment length $\ell_{\text{a}}$. A schematic illustration is presented in Fig. 6c. Ignoring the noise terms $\bm{\eta}_{i}^{\text{t}}$ and $\bm{\eta}_{i}^{\text{r}}$, we substitute the velocity in the second equation using the first equation, thereby obtaining the following expression for the equations of motion
@@
which is equivalent to Eq. (B.3), except for the noise terms.
B.4 Programmed robotic metamaterials
Let us now look at mechanical metamaterials made of robots whose interactions are programmed to break reciprocity  [17, 34]. In such materials, nonreciprocal interactions give rise to spatially asymmetric standing waves and unidirectionally amplified propagating waves, realizing the mechanical analogue of the non-Hermitian skin effect [59, 88]. Specifically, the robots are programmed to turn more (less) depending on how their left (right) neighbors turn, as schematized in Fig. 6d. The dynamics can be mapped onto the following nonreciprocal discrete wave equation, which arises from a mass-spring model in which the forces exerted by the springs on the left and right sides differ [17]:
@@ B.4 Programmed robotic metamaterials
Here $u_{i}$ denotes the displacement of robot $i$; $k\left(1+\varepsilon\right)$ and $k\left(1-\varepsilon\right)$ represent the stiffnesses of coupling springs from left to right and from right to left, respectively, with the dimensionless quantity $\varepsilon$ quantifying nonreciprocity. For convenience, we have set the mass to unity.
@@
We observe that the kinetic energy of the auxiliary particles is negative. Such a negative kinetic energy term in the auxiliary system is necessary to enforce our constraints; a more detailed explanation will be provided in Sec. C.2.
@@
B.5 Feedback-controlled active particles with vision-cone interactions
@@
where $V_{i}^{\alpha}$ denotes the vision cone of angle $\alpha$ around the orientation of particle $i$ (Fig. 6e). Respectively, $r_{ij}$ is the distance between particles $i$ and $j$, so that the perception metric assigns less weight to more distant particles. If a given particle has a perception metric $P_{i}>P^{*}$, it is made to propel at speed $v_{0}$, which is achieved by shining light on it through the laser scanning system. Otherwise, the particle remains passive. Thus, particle motion can be modeled as
@@
which is the same as Eq. (B.5), except for the noise terms.
B.6 Sedimenting particles with hydrodynamic interactions
Sedimenting particles provide a paradigmatic example of a system with velocity-dependent nonreciprocal interactions due to hydrodynamics. When particles settle through a viscous fluid under gravity, they generate flow fields that affect the motion of neighboring particles. In general, the flow generated by particle $j$ impacts particle $i$ differently than the flow generated by particle $i$ impacts particle $j$ (Fig. 6f). Therefore, the resulting interactions are intrinsically nonreciprocal. This nonreciprocity arises not from the particles themselves, but from the asymmetry of the hydrodynamic coupling through the surrounding fluid.
Consider two spherical particles sedimenting under gravity in a Stokes flow regime, where inertial effects are negligible. The equation of motion for particle $i$ reads as
$\gamma\dot{\mathbf{r}}_{i}=m_{i}\mathbf{g}+\sum_{j\neq i}\mathbf{F}_{j\to i}^{\text{hydro}}+\bm{\eta}_{i},$ (112)
where $m_{i}$ is the particle mass, $\mathbf{g}$ is gravitational acceleration, $\gamma$ is the Stokes drag coefficient (proportional to the fluid viscosity), $\mathbf{F}_{j\to i}^{\text{hydro}}$ denotes the hydrodynamic force exerted on particle $i$ by the flow generated by particle $j$, and $\bm{\eta}_{i}$ is the noise term.
The hydrodynamic interaction force can be expressed in terms of the Oseen tensor $\mathbf{G}$ (also called the Stokeslet):
$\mathbf{F}_{j\to i}^{\text{hydro}}=\gamma\mathbf{G}(\mathbf{r}_{ij})\cdot m_{j}\mathbf{g},$ (113)
where $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$ is the inter-particle separation, and the Oseen tensor is given by
$\mathbf{G}(\mathbf{r})=\frac{1}{8\pi\eta r}\left(\mathbf{I}+\frac{\mathbf{r}\otimes\mathbf{r}}{r^{2}}\right),$ (114)
with $\eta$ the dynamic viscosity of the fluid, $\mathbf{I}$ the identity matrix, and $\mathbf{r}\otimes\mathbf{r}]_{\alpha\beta}=r_{\alpha}r_{\beta}$. The symmetry of this tensor implies that $\mathbf{G}(\mathbf{r}_{ij})=\mathbf{G}(\mathbf{r}_{ji})$, reflecting the directional dependence of hydrodynamic coupling on the separation vector. Therefore, $\mathbf{F}_{j\to i}^{\text{hydro}}$ and $\mathbf{F}_{i\to j}^{\text{hydro}}$ point in the same direction since the masses of particles are all positive. Hence, the interaction forces are not equal and opposite; the interactions are nonreciprocal. Crucially, the nonreciprocal interactions in this system are velocity-independent: they depend only on particle positions through the inter-particle distance.
We can now introduce auxiliary particles with position denoted by $\mathbf{x}_{i}$. The Hamiltonian embedding is given by
$H=-\sum_{i}m_{i}\mathbf{g}\cdot\mathbf{r}_{i}-\sum_{\langle ij\rangle}\mathbf{F}_{j\to i}^{\text{hydro}}\cdot(\mathbf{x}_{i}+\mathbf{r}_{i}).$ (115)
The corresponding equation of motion from this Hamiltonian description reads as
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}=-\frac{\partial H}{\partial\mathbf{r}_{i}}$ $=m_{i}\mathbf{g}+\sum_{j\in\langle ij\rangle}\mathbf{F}_{j\to i}^{\text{hydro}}+\sum_{j\in\langle ij\rangle}\frac{\partial}{\partial\mathbf{r}_{i}}\mathbf{F}_{j\to i}^{\text{hydro}}\cdot(\mathbf{x}_{i}+\mathbf{r}_{i}).$ (116)
Moreover, the constraint is
$\mathbf{x}_{i}(t)+\mathbf{r}_{i}(t)=0.$ (117)
Applying the constraints to the equations of motion yields
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=m_{i}\mathbf{g}+\sum_{j\in\langle ij\rangle}\mathbf{F}_{j\to i}^{\text{hydro}}\;,$ (118)
which is equivalent to Eq. (112), up to the noise term.
Appendix C Symplectic structure of the Hamiltonian embedding: Implementing the constraint for interactions with overdamped and undamped nonreciprocal dynamics
In Sec. A, we introduced Hamiltonian embeddings and outlined the appropriate constraints for several nonreciprocal systems, both for damped and undamped dynamics in Sec. B. Here, we elaborate on these two cases separately, discussing in detail the implementation of the constraint and the corresponding symplectic structure. Although the discussion below is general, for concreteness we shall illustrate the procedure using the vision-cone XY spins as an example, for which the interaction $J_{ij}(\theta_{i})$ takes the form
@@ B.6 Sedimenting particles with hydrodynamic interactions
C.1 Overdamped nonreciprocal systems
@@
Using Eq. (C.1), the equation of motion of the constraint is given by
@@
$\sin\left(\theta_{i}-\theta_{j}\right)+\sin\left(\theta_{i}-\varphi_{j}\right)$ $=$ $0.$ (126)
@@
C.2 Undamped nonreciprocal systems
@@
Using Eq. (C.2), the equations of motion of the constraints, $G^{(1)}_{i}$ and $G^{(2)}_{i}$, are given by
@@
$\sin\left(\theta_{i}-\theta_{j}\right)+\sin\left(\varphi_{i}-\theta_{j}\right)$ $=$ $0.$ (137)
@@
The occurrence of negative kinetic energy for the auxiliary degrees of freedom can be explained as follows: Consider first the free system where the degrees of freedom are noninteracting, $\mathcal{U}=0$, and both terms in the kinetic energy are positive; then, by conservation of angular momentum $L^{\varphi}_{j}+L^{\theta}_{j}=0$, the $\theta$ and $\varphi$ degrees of freedom will rotate opposite to one another, thus violating the constraint that pins them to move together (i.e., in the same direction). To enforce the latter without over-constraining the dynamics of the resulting joint motion, a negative kinetic energy is required. Returning to the Hamiltonian (C.2), more formally, as discussed above, we need to construct a constraint as $\theta_{i}(t)=\varphi_{i}(t)+\pi$. Taking two time derivatives, we have $\ddot{\theta}_{i}(t)=\ddot{\varphi}_{i}(t)$. If the interactions between the original and the auxiliary degrees of freedom are defined as in Eq. (83), the torques experienced by the original spin and the auxiliary spin will be oppositely directed under the imposed constraint, as given by $\sum_{j\in\langle ij\rangle}J_{ij}\sin(\theta_{i}-\theta_{j})=-\big[-\sum_{j\in\langle ij\rangle}J_{ij}\sin(\varphi_{i}-\theta_{j})\big]$. To maintain the opposite sign in the EOM of the angular momenta, the kinetic energy for the auxiliary degree of freedom has a negative sign, $\frac{1}{2}\dot{\varphi}_{i}^{2}$. Such negative kinetic energy terms for the auxiliary degree of freedom have been discussed in Ref. [32] where they arise as a result of reformulating the action principle as an initial- (rather than a boundary-) value problem. Similarly, in the programmed robotic metamaterials example from Sec. B.4, we used the kinetic term $\frac{1}{2}\sum_{i}\left(p_{i}^{u}\right)^{2}-\left(p_{i}^{w}\right)^{2}$ to construct the Hamiltonian. Such anisotropic kinetic energy terms have also been encountered before in the description of azimuthal curl forces [12].
@@
Appendix D Equivalence between nonreciprocal Langevin dynamics and Glauber Monte-Carlo dynamics
This section demonstrates that the steady-state and non-stationary density distributions resulting from the original nonreciprocal Langevin dynamics and from the constrained Glauber dynamics, based on the Hamiltonian that we propose, are identical. In Sec. D.1, we first derive a Fokker-Planck-like equation for the phase-space density distribution obtained from Langevin dynamics; then, in Sec. D.2, we prove that it is equivalent to the corresponding distribution obtained using constrained Glauber dynamics.
@@ Appendix D Equivalence between nonreciprocal Langevin dynamics and Glauber Monte-Carlo dynamics
although the procedure is general and applies to other models (see Sec. B).
Before presenting the detailed derivations, we briefly clarify the notions of detailed balance and global balance [47] for Markov dynamics. Let $\mathcal{C}$ denote a microstate (configuration), e.g., the set of spin angles $\{\theta_{i},\varphi_{i}\}$ of the Hamiltonian embedding in Eq. (141), and let $W(\mathcal{C}\to\mathcal{C}^{\prime})$ be the transition rate of the Markov dynamics from $\mathcal{C}$ to $\mathcal{C}^{\prime}$.
Detailed balance requires the probability current to vanish pairwise:
$\rho(\mathcal{C})\,W(\mathcal{C}\to\mathcal{C}^{\prime})=\rho(\mathcal{C}^{\prime})\,W(\mathcal{C}^{\prime}\to\mathcal{C})\qquad\forall\,\mathcal{C},\mathcal{C}^{\prime}.$ (142)
This condition implies the absence of steady-state probability currents and is characteristic of equilibrium dynamics.
Global balance is weaker: it only requires that, for every microstate, the total incoming probability flux equals the total outgoing flux,
$\sum_{\mathcal{C}^{\prime}}\rho(\mathcal{C}^{\prime})\,W(\mathcal{C}^{\prime}\to\mathcal{C})=\rho(\mathcal{C})\sum_{\mathcal{C}^{\prime}}W(\mathcal{C}\to\mathcal{C}^{\prime}),\qquad\forall\,\mathcal{C},$ (143)
thereby ensuring stationarity while allowing for nonzero circulating currents. Since nonreciprocal interactions generically drive the system out of equilibrium, detailed balance is not expected to hold; accordingly, in the following, we take global balance to be the appropriate steady-state condition. We will state the global-balance condition in its most general form and characterize the class of stationary distributions consistent with it in our setting. At the same time, in specific models a stronger condition (intermediate between global and detailed balance) may hold; for instance, in Eq. (143) the sum over $\mathcal{C}^{\prime}$ may effectively be restricted to the subset of configurations that are dynamically connected to $\mathcal{C}$. Also, note that detailed balance implies global balance, but not vice versa.
Finally, if global balance in Eq. (143) is violated, the density distribution is non-stationary. In that case, the time evolution of the density is governed by the master equation
$\partial_{t}\rho(\mathcal{C})=\sum_{\mathcal{C}^{\prime}}\rho(\mathcal{C}^{\prime})\,W(\mathcal{C}^{\prime}\to\mathcal{C})-\rho(\mathcal{C})\sum_{\mathcal{C}^{\prime}}W(\mathcal{C}\to\mathcal{C}^{\prime}),\qquad\forall\,\mathcal{C}.$ (144)
Although we do not use this relation directly in our proof, it can be recovered naturally from the derivations below.
D.1 Langevin steady-state distribution